OpenStudy (anonymous):

sec^6x(secxtanx)-sec^4x(secxtanx)=sec^5xtan^3x verify

OpenStudy (alekos):

i have the method but have you tried this?

OpenStudy (anonymous):

tried what?

OpenStudy (alekos):

if we look at the LHS we have secxtanx present in both terms, so this can be factored out comes to (sec^6x - sec^4x)(secxtanx) what do you think the next step will be?

OpenStudy (anonymous):

oh well it becomes sec^2x which becomes 1+tan^2x

OpenStudy (anonymous):

actually nevermind that doesn't work I don't know what to do next

OpenStudy (alekos):

next step is (sec^4x.sec^2x - sec^4x)(secx.tanx) you follow so far

OpenStudy (anonymous):

so you just broke apart the sec^6x

OpenStudy (alekos):

yep you got it, now for the next step [sec^4x(sec^2x -1)](secxtanx) happy with that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

thank you! can I ask another

OpenStudy (alekos):

we know that sec^2x = 1+tan^2x so we have (sec^4x.tan^2x)(secxtanx) what do you think we do now?

OpenStudy (anonymous):

multiply and your done

OpenStudy (alekos):

that's it!! and it equals the RHS

OpenStudy (anonymous):

six(1-2cos^2x+cos^4x)=sin^5x

OpenStudy (alekos):

ok give me a few minutes

OpenStudy (anonymous):

got it

OpenStudy (alekos):

whats the six doing right at the front?

OpenStudy (anonymous):

there is no six?

OpenStudy (alekos):

This is what you wrote six(1-2cos^2x+cos^4x)=sin^5x

OpenStudy (anonymous):

yeah I don't see a six?

OpenStudy (anonymous):

oh the word six its sinx

OpenStudy (alekos):

that's what i thought it might be

OpenStudy (anonymous):

I factored the cos^4x and got no where

OpenStudy (alekos):

what did you factor it to?

OpenStudy (anonymous):

(1-cos^2x)^2

OpenStudy (alekos):

oh i see should be (cos^2x -1)^2

OpenStudy (anonymous):

got it that doesn't simplify tho to something else

OpenStudy (alekos):

yes it does! we know that cos^2x + sin^2x =1 so what do you think that cos^2x - 1 will equal?

OpenStudy (anonymous):

-sin^2x but how does that equal positive sin on the other side

OpenStudy (anonymous):

oh its squared I got it

OpenStudy (alekos):

well done

OpenStudy (alekos):

it becomes sin^4x.sinx which is equal to the RHS

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

4tan^4x+tan^2x-3=sec^2x(4tan^2x-3)

OpenStudy (alekos):

ok, let's do this together, give me a min or two

OpenStudy (anonymous):

I have this and one more then im done

OpenStudy (alekos):

OK standby

OpenStudy (alekos):

this time we look at the RHS.... we know that sec^2x = tan^2x + 1 so we have... (tan^2x +1)(4tan^2x - 3) now what's next?

OpenStudy (anonymous):

really you just multiply and that's it?

OpenStudy (anonymous):

why couldn't I notice that before

OpenStudy (alekos):

because these problems require practice, a keen eye and lots of thinking. i think you're on your way to achieving all three

OpenStudy (alekos):

multiply this out and you get the LHS

OpenStudy (anonymous):

(1+sec(-x))/(sin(-x)+tan(-x))=-cscx

OpenStudy (alekos):

ok, standby

OpenStudy (anonymous):

yes

OpenStudy (alekos):

ok here we go, this one is a bit more tricky...... we know that sec(-x) = secx and sin(-x) = -sinx and tan(-x) = -tanx so looking at the LHS we now have......... (1 + secx)/-(sinx +tanx) happy with that? what do you think we do now?

OpenStudy (anonymous):

why does the negative for sinx go outside the( )

OpenStudy (alekos):

because sin(-x) + tan(-x) = -sinx - tanx = -(sinx + tanx)

OpenStudy (anonymous):

this is true yes

OpenStudy (alekos):

ok, got any ideas for the next step?

OpenStudy (anonymous):

this is where I have no idea

OpenStudy (alekos):

ok well whats another way of expressing secx? and tanx?

OpenStudy (anonymous):

1/cosx sinx/cosx

OpenStudy (anonymous):

like I know you can swap those but then it gets ugly

OpenStudy (alekos):

yes you're right and another way of expressing tanx is sinx.secx you ok with that?

OpenStudy (anonymous):

yeah

OpenStudy (alekos):

ok so we substitute sinx.secx for tan x in the bottom line and we get (1 + secx)/-(sinx + sinx.secx) can you see where i'm going?

OpenStudy (anonymous):

no it still looks like I have no idea where to go after that

OpenStudy (alekos):

ok look at it carefully the bottom line can now be factored out with sinx can you see?

OpenStudy (anonymous):

well yeah but how does that help with the top

OpenStudy (alekos):

because then you'll have (1 + secx) on the top and bottom which cancels out!

OpenStudy (anonymous):

I see I was like where are you going with this...but see I would never realize that on my own

OpenStudy (alekos):

so we're left with -1/sinx which is equal to the RHS

OpenStudy (alekos):

if you look at things hard enough and do a bit of trial and error with heaps of logical thinking then you can do this too!

OpenStudy (alekos):

always use the basic trig identities to work these questions out

OpenStudy (alekos):

was that the last one?

OpenStudy (anonymous):

yes so your saying when you can switch from cos^2x to 1-sin^2x to always try and do that?

OpenStudy (alekos):

yes that's one of the many trig identities, there's heaps more. just look at your text books or the internet, but that's the basic idea. trial and error until you work it out, then you"ll just get better at it

OpenStudy (alekos):

see you next time